OpenStudy (anonymous):

Which pairs are solutions to the equation? 3xy + 9 = 33 A. (1, 8) and (1, 11) B. (1, 11) and (2, 6) C. (1, 8) and (8, 1) D. (8, 1) and (2, 6)

OpenStudy (hesan):

How it is possible to get value of two unknown by single equation???

OpenStudy (hesan):

@theoryofmars Are u getting my point??

OpenStudy (texaschic101):

I think it is C. (1,8) 3xy + 9 = 33 3(1)(8) + 9 = 33 24 + 9 = 33 33 = 33 (correct) (8,1) 3xy + 9 = 33 3(8)(1) + 9 = 33 24 + 9 = 33 33 = 33 (correct)